OpenStudy (anonymous):

(Cauchy Mean Value Theorem) Suppose the function \[f:(-1,1)\to\mathbb{R}\] has n derivatives and that its nth derivative \[f^{(n)}:(-1, 1)\to\mathbb{R} \]is bounded. Also assume that \[f(0) = f'(0) = \dotso = f^{(n-1)}(0) = 0.\] Prove \[\exists M>0 : |f(x)| \leq M|x|^{n} \\\forall x \in (-1, 1).\] Probably more effort to LaTeX that than figure it out, oh well...

OpenStudy (anonymous):

i have an idea not sure if it is correct

OpenStudy (anonymous):

no it isn't let me try again

OpenStudy (anonymous):

The solution almost definitely involves a watered down version of the Lagrange Remainder Theorem that they give before defining Taylor polynomials (since Lagrange remainders are ultimately derived by Cauchy Mean Value), which I can type up in a second...

OpenStudy (anonymous):

Theorem: Let I be an open interval and n be a natural number and suppose that the function \[f:I\to\mathbb{R}\] has n derivatives. Suppose also that at \[x_0 \in I, f^{(k)}(x_0) = 0; 0\leq k\leq n-1.\]Then \[\forall x\neq x_0, x\in I \]There exists a z strictly between x, x0 such that \[f(x) = \frac{f^{(n)}(z)}{n!}(x-x_0)^n. \]

OpenStudy (anonymous):

So here we're given just such an f that that final form can be used, I guess we can set \[M = \frac{f^{(n)}(z)}{n!}\]And then monkey with it somehow...

OpenStudy (anonymous):

i am no help, sorry

OpenStudy (anonymous):

No worries!

OpenStudy (anonymous):

Yeah I was right. Given that our function is n-times differentiable on an open interval etc etc and given that we have \[f^{(k)}(0) = 0, 0\leq k \leq n-1 \] it's permissable to set x0 = 0 and the theorem becomes \[f(x) = \frac{f^{(n)}(z)}{n!}(x)^n, 0<|z|\leq 1 \]Thus, \[|f(x)| = \bigg|\frac{f^{(n)}(z)}{n!}x^n\bigg| = \bigg|\frac{f^{(n)}(z)}{n!}\bigg||x|^n \]And therefore, \[|f(x) \leq M|x|^n \]for \[M\geq \bigg|\frac{f^{(n)}(z)}{n!}\bigg|\quad\Box \]